Projection Admin 1/26
To do this evening/ was completed: Possessor keys open up tomorrow! Can get tech'd first thing in the morn.
Presentation Issues:
Preshow Clips/Playlist: Finalized the playlist(s) for Hank's Saloon. Tested spots/mics. See more notes in Special Events. Leon's drive is on the table in 4/6, in its box, all ready to get handed back to him. Spots setting for centre spots in T2 is 301. LH
Double checked Road Games. LH
Trailer Updates:
Special Events/ AV set up: DJ is coming in at 7:30 for Hank's Saloon afters. I did a preset with the xlr cables into the dj inputs, the dual rca to xlr is on the table though he will bring his own cabling for this I figured it would be best to have it out anyway, along with a powerstrip and extension cord. It should be a super simpe setup. Showed Jack, who has the dj key to turn everything on. Lauren and Ariel reached out to say they'd take the reins on setting him up, but we'll make sure to check in. I checked in with them this evening already. - LH
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): DCP'd the last short for NoBudge, Goodnight, which is now copying over to the share folder with all the other shorts. Please ingested them once they've copied over. Also, we received the music for the performances, which we can now DCP. I downloaded these to the desktop. LH
Outgoing/Shipping (content): IMPORTANT - I am copying Road Games over to the preshows/trailers drive so it doesn't have to stay on the share folder. We will soon receive a blank drive from The Academy to put the DCP on, then we'll send it to them. Please DO NOT delete this feature from the preshows/trailers drive. LH
Brough The Birds down to the black cart, which will get picked up tomorrow via FedEx Ground. This print came with return labels (NBC Universal). I put the tracking in the spreadsheet and scheduled the pickup. LH
Print/video Inspection:
Asset tracking doc updated/checked?: Yes
Maintenance/cleaning: Adjusted lens files in T2. Will alerted me that the rear surround speakers were not working in T1 when running the dolby 7.1 test. Few things:
Trailer Updates:
Special Events/ AV set up: DJ is coming in at 7:30 for Hank's Saloon afters. I did a preset with the xlr cables into the dj inputs, the dual rca to xlr is on the table though he will bring his own cabling for this I figured it would be best to have it out anyway, along with a powerstrip and extension cord. It should be a super simpe setup. Showed Jack, who has the dj key to turn everything on. Lauren and Ariel reached out to say they'd take the reins on setting him up, but we'll make sure to check in. I checked in with them this evening already. - LH
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): DCP'd the last short for NoBudge, Goodnight, which is now copying over to the share folder with all the other shorts. Please ingested them once they've copied over. Also, we received the music for the performances, which we can now DCP. I downloaded these to the desktop. LH
Outgoing/Shipping (content): IMPORTANT - I am copying Road Games over to the preshows/trailers drive so it doesn't have to stay on the share folder. We will soon receive a blank drive from The Academy to put the DCP on, then we'll send it to them. Please DO NOT delete this feature from the preshows/trailers drive. LH
Brough The Birds down to the black cart, which will get picked up tomorrow via FedEx Ground. This print came with return labels (NBC Universal). I put the tracking in the spreadsheet and scheduled the pickup. LH
Print/video Inspection:
Asset tracking doc updated/checked?: Yes
Maintenance/cleaning: Adjusted lens files in T2. Will alerted me that the rear surround speakers were not working in T1 when running the dolby 7.1 test. Few things:
1. I could not connect to any cp750 using the dolby app, which made it difficult to check channel activity more closely.
2. It was, in fact, the right sound test for all channels.
3. Amplifiers were/are on, but no movement during their (the rear surrounds) turn.
4. Powercycled the cp750 and it's ethernet cable, tried a different port on the switch too, to no avail.
5. Everything else is working just fine. These speakers might need to be replaced and should absolutely get checked out tomorrow. LH
Cue sheets updated/checked (every shift): Updated and corrected all SEs for today. LH
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Cue sheets updated/checked (every shift): Updated and corrected all SEs for today. LH
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Comments
Post a Comment